OpenStudy (anonymous):

A car travels 17 miles in 1 gallon of gas. Express y in terms of x if the car uses y gallons to travel x miles.

OpenStudy (radar):

From what is given, how many miles per gallon would be the rating for this car?

OpenStudy (anonymous):

17 miles per gallon

OpenStudy (radar):

Y is to be gallons, x is miles. If you drove 34 miles, x = 34 and y would be 2 gallons. The problem wants you to set up a function f(y) that would solve the number of gallons for a given number of miles. f(y) = x/17 or f(y) = (1/17) x Now to test it How many gallons of gas is used to travel 85 miles? f(y) = 85/17 = 5 gal

OpenStudy (radar):

You see the 17 is a constant, y is the dependent variable, the x is the independent variable.

OpenStudy (anonymous):

Great! Thank you so much. Can I work it out as a proportion? I can do something like 17/1 = x/y

OpenStudy (anonymous):

so I get the same answer like yours y = x/17

OpenStudy (radar):

Thats aother way to look at it. 17:1 :: 85:y 85 = 17y y=5 gal Neat way to look at it. Good luck with your studies.

OpenStudy (anonymous):

Thanks a ton. You've been so helpful